OpenStudy (anonymous):

What is the slope of the line between (3, −4) and (−2, 1)? 1 2 −2 −1

OpenStudy (zzr0ck3r):

The slope \(m\) of the line between \((x_1,y_1)\) and \((x_2,y_2)\) is given by \[m=\dfrac{y_1-y_2}{x_1-x_2}\]

OpenStudy (anonymous):

So it'd be -1?

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

Thanks!

OpenStudy (zzr0ck3r):

np

OpenStudy (zzr0ck3r):

note that \[\dfrac{y_1-y_2}{x_1-x_2}=\dfrac{-(y_2-y_1)}{-(x_2-x_1)}=\dfrac{y_2-y_1}{x_2-x_1}\] So you can do it in either order, just make sure that the \(y_2\) goes with the \(x_2\) and the \(y_1\) goes with the \(x_1\).